OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin2x + sin x = 0

OpenStudy (agent0smith):

Use the fact that sin2x = 2sinxcosx first. Then factor out sinx. Nice fake pic btw :P

OpenStudy (luigi0210):

btw, is it \(sin(2x)\) or \(sin^2x\).. big difference

OpenStudy (anonymous):

lol @agent0smith

OpenStudy (anonymous):

Thanks. It's the best way to attract pervs and trick them into doing my work. Dueche.@agent0smith Nice name by the way. Agent Smith. It's sin^2x. Sorry about that @Luigi0210

OpenStudy (agent0smith):

LOL thanks. Good luck getting help, calling people... whatever that is. Looks german.

OpenStudy (agent0smith):

If it's sin^2 x then just factor out sinx right now.

OpenStudy (agent0smith):

Right click pic => google images Or just use a less completely obviously fake pic than that one, wouldn't set off red flags in the first place.

OpenStudy (anonymous):

a. x = 0, π, 4π/3, 5π/3 b. x = 0, π, π/3, 2π/3 c. x = 0, π, 3π/2

OpenStudy (anonymous):

those are my options

OpenStudy (agent0smith):

So... start by factoring out sinx...

OpenStudy (shadowlegendx):

@girlfromafrica is it possible if I can marry you?

OpenStudy (anonymous):

It's a British Pakistanian

OpenStudy (anonymous):

And I don't care if people know my pic is fake. I don't want the internet world to see my face. Darling @ShadowLegendX This face is fake. You wouldn't want it.

OpenStudy (anonymous):

@ShadowLegendX she could be a dude lol

OpenStudy (elisaneedshelp):

@A_Plus_Forever

OpenStudy (elisaneedshelp):

i replied on chat

OpenStudy (anonymous):

Lol I'm not a dude.

OpenStudy (anonymous):

and how would we know that? lol

OpenStudy (shadowlegendx):

Fine...I will go marry the 60year old german man in English chat @adrynicoleb

OpenStudy (agent0smith):

sin^2 x + sin x = 0 sooo factor out a sinx and what do you get?? @girlfromafrica

OpenStudy (anonymous):

Lol I don't care what you thing. 0. I get 0. @agent0smith

OpenStudy (anonymous):

Think

OpenStudy (agent0smith):

Show me what you get. It can't just be 0.

OpenStudy (anonymous):

1^2 + 0?

OpenStudy (adrynicoleb):

LOL I'm cool with that ;)

OpenStudy (anonymous):

If I change my pic to hot girl with big boobs, would anyone answer my question i just posted? :3

OpenStudy (anonymous):

Haha they would @sourwing

OpenStudy (agent0smith):

@girlfromafrica not quite sin^2 x + sin x = 0 factor out sinx then you should get sinx(sinx + 1) = 0

OpenStudy (anonymous):

All sorts of pervs would message you being like: "heyyy babe. you got snapchat?" LOL @sourwing

OpenStudy (anonymous):

No,....................... it has to be a big retriceAND big boobs

OpenStudy (anonymous):

*big butt

OpenStudy (anonymous):

Ohhhh okay. @agent0smith What's the next step?

OpenStudy (elisaneedshelp):

yo i cant reply much over pm cuz my boyfriend will see might see

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@ElisaNeedsHelp Then stop talking to random strangers

OpenStudy (anonymous):

She's not lol

OpenStudy (agent0smith):

@girlfromafrica then you set each equal to zero: sinx(sinx + 1) = 0 sinx = 0 sinx + 1 = 0 You should hopefully be able to solve from here.

OpenStudy (anonymous):

You guys know each other? Ooooo y'all are playing with fire. @A_Plus_Forever @ElisaNeedsHelp

OpenStudy (anonymous):

She doesent want her bf getting mad that she's talking to another boy but she's just talking but he'll still get mad

OpenStudy (anonymous):

Guys. I need help getting the answer. Please. Options: a. x = 0, π, 4π/3, 5π/3 b. x = 0, π, π/3, 2π/3 c. x = 0, π, 3π/2

OpenStudy (anonymous):

Boys get like that.

OpenStudy (elisaneedshelp):

cuz its not just you its so many other messages i get and other ppl are really weird

OpenStudy (anonymous):

Lol no kidding. @ElisaNeedsHelp

OpenStudy (anonymous):

Puberty.

OpenStudy (agent0smith):

@girlfromafrica attempt to solve what I gave you. sinx = 0 sinx + 1 = 0

OpenStudy (elisaneedshelp):

so i have many weird boy messages Your one of the normal ones lol

OpenStudy (anonymous):

preech bruttha @A_Plus_Forever

OpenStudy (anonymous):

How? All I get is 0. 0 0 0 0 0. That's all I get. @agent0smith

OpenStudy (anonymous):

Who wants long D?

OpenStudy (shadowlegendx):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}D\color{#0000ff}a\color{#2400ff}t\color{#6D00ff}i\color{#B600ff}n\color{#FF00ff}g\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)

OpenStudy (anonymous):

lol jkjkjkjkjk

OpenStudy (anonymous):

Stooooop. @A_Plus_Forever

OpenStudy (shadowlegendx):

Find your dream girl ;)

OpenStudy (agent0smith):

@girlfromafrica what do you mean you just get 0. Do you mean x=0? sinx = 0 sinx + 1 = 0 ^where does sinx=0?? ^you've made no attempt to solve this one.

OpenStudy (elisaneedshelp):

lol @A_Plus_Forever dont get like the other guys plz you have a good shot with somebody

OpenStudy (anonymous):

How can I solve something that = 0? @agent0smith

OpenStudy (elisaneedshelp):

Lol Opendating

OpenStudy (agent0smith):

@girlfromafrica do you have any idea how to solve trigonometric equations? What value(s) of x, will make sinx=0? sinx + 1 = 0, for this one, first you have to subtract 1 from both sides. sinx = -1 What values of x will make sinx = -1? Use a unit circle. Or something. Do not just say "i get 0".

OpenStudy (anonymous):

Lol yeah right I cant get a girlfriend for shizballs

OpenStudy (elisaneedshelp):

Lol im sure thats not true you proved me your a really good person :) and the words you said show you are worth it for somebody

OpenStudy (elisaneedshelp):

You will get a gf *wink wink*

OpenStudy (anonymous):

So I'm getting x = 0, π/3, π, 5π/3 as the final answer @agent0smith

OpenStudy (agent0smith):

Did you guess that as your answer? Because you answered zero questions from my last post.

OpenStudy (anonymous):

No I didn't guess. I'm just trying to finish my hw while doing this

OpenStudy (johnweldon1993):



OpenStudy (agent0smith):

Then if you didn't guess, explain HOW you got that answer.

OpenStudy (johnweldon1993):

So to solve sin(x) = 0 Looks like that happens at 2 points on that unit circle....which 2 points are those?

OpenStudy (anonymous):

The one on the wayleft and the one on the way right @johnweldon1993

OpenStudy (anonymous):

Rewrite sin(2x) as 2sin(x)cos(x): sin(x) + 2sin(x)cos(x) = 0 Factor out sin(x): sin(x)[1 + 2cos(x)] = 0 Set each factor to equal zero: sin(x) = 0 x = 0, π 1 + 2cos(x) = 0 2cos(x) = -1 cos(x) = -1/2 x = π/3, 5π/3 @agent0smith

OpenStudy (agent0smith):

That's not the same problem as we've been doing... you pointed out it was sin^2 x not sin2x.

OpenStudy (anonymous):

Okay so....I was wrong. @agent0smith

OpenStudy (anonymous):

It's actually sin2x

OpenStudy (agent0smith):

If you're able to solve that one, then you're able to solve this one.

OpenStudy (johnweldon1993):

@girlfromafrica correct... "The one on the way left" is resembled by \(\pi\) and the one of the far right resembles \(2 \pi\) which is also = to 0... o far we have 2 points...\(\pi\) and 0 Now...onto the second equation @agent0smith has given you above... sin(x) + 1 = 0 subtract 1 from both sides of the equation sin(x) = -1 Where does sin(x) = -1 on that unit circle?

OpenStudy (agent0smith):

Oh wow.

OpenStudy (johnweldon1993):

@girlfromafrica Wait...it's sin(2x)?? -_-

OpenStudy (anonymous):

I'M SORRY

OpenStudy (anonymous):

Am I working it out right though?

OpenStudy (anonymous):

I don't think my answer is right.

OpenStudy (agent0smith):

Easy way to check: plug the options into sin2x + sin x = 0 and see which make the equation true. I don't understand how you went from saying "I get 0" to... writing out a complete solution...

OpenStudy (agent0smith):

https://answers.yahoo.com/question/index?qid=20140124114101AAriB8J one of those looks awfully familiar...